OpenStudy (anonymous):

Write down the Pythagorean Identity sin^2 x +cos^2 x = 1. Divide each side of the equation by sin^2 x and describe the simplifications that can be done. I know that the Pythagorean identity for sin2 x + cos2 x = 1 is sin2u + cos2u = 1 but I don't know where to go from there.

OpenStudy (anonymous):

divide each term by \(\sin^2(x)\) and see what you get first step is \[\frac{\sin^2(x)}{\sin^2(x)}+\frac{\cos^2(x)}{\sin^2(x)}=\frac{1}{\sin^2(x)}\]

OpenStudy (anonymous):

then rewrite each term in terms of a trig function well, except for the first term which is evidently \(1\)

OpenStudy (anonymous):

I get the first step. But what is the format for a trig function?

OpenStudy (anonymous):

@mathmath333 @Algorithmic @gorv @Eagleeye13 @helpout3 @hugsnkisses @iGreen

OpenStudy (triciaal):

@sunshine201 are you suppose to prove sin^2 x +cos^2 x = sin2u + cos2u ?

OpenStudy (anonymous):

I read the problem wrong. No. can you rewrite each term in terms of a trig function?

OpenStudy (triciaal):